The problem is one number is 2 more then 5 times another number their sum is 24. what are the number. I need the steps on how to solve the problem
Answer by ankor@dixie-net.com(6693)

Let x= " a number"
Let y = "another number"
:
Write an equation for what it says (the word "is" usually means =:
:
"one number is 2 more then 5 times another number" their sum is 24."
x = 5y + 2
:
"their sum is 24"
x + y = 24
x = (24-y); this form for substitution
:
what are the numbers?
:
Substitute (24-y) for x in the 1st equation
24 - y = 5y + 2
:
24 - 2 = 5y + y
:
22 = 6y
y = =
y = 3
:
Find x
x = 5y + 2
x = 5() + 2
x = + 2
x = 18 + 2
x = 20
:
:
Check by finding the sum:
20 + 3 = 24